A 10kg mass has a ke of 125 j what is the velocity



1
Expert's answer
2022-02-08T08:16:33-0500

Given:

"m=10\:\rm kg"

"E_k=125\:\rm J"


The rinetic energy of an object

"E_k=\frac{mv^2}{2}"

Hence, the velocity

"v=\sqrt{\frac{2E_k}{m}}=\sqrt{\frac{2*125}{10}}=5\:\rm m/s"